OpenStudy (iwanttogotostanford):

QUICK ALGEBRA QUESTION< NEED HELP_Expand the 5th root of x into an exponent with a fraction, 1/5 In the exponent of x, multiply 7 times this value In the exponent of x, Multiply the result times 3

OpenStudy (anonymous):

what O.o (, )_( ')

OpenStudy (iwanttogotostanford):

SORRY NOT THAT : Describe how to transform the quantity of the fifth root of x to the seventh power, to the third powerinto an expression with a rational exponent. Make sure you respond with complete sentences.@Dyron_Libaunask

OpenStudy (iwanttogotostanford):

\[(\sqrt[5]x{^7})^3\]

OpenStudy (anonymous):

@THEHELPER123

OpenStudy (iwanttogotostanford):

Thats the equation

OpenStudy (anonymous):

he2 better at explenatiion2 than ii am

OpenStudy (iwanttogotostanford):

if he doesn't show up can you still just explain step by step please?

OpenStudy (anonymous):

dumb iit down, what do you need for an awncer?

OpenStudy (iwanttogotostanford):

i need basically a short paragraph on explaining how to solve this problem

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

pemdas (5√x7)3 ... XC ii 2uck at explenatiion2

OpenStudy (anonymous):

2orry, ii can get you an awncer two iit but iit2 allmental math and ii dont do well whiith 2howiing work

OpenStudy (anonymous):

wiith*

OpenStudy (iwanttogotostanford):

ok

OpenStudy (anonymous):

ive actually never learned about this yet

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

wow O.O

OpenStudy (anonymous):

ik im not good a exponets i never heard of the 5th root

OpenStudy (iwanttogotostanford):

its hard i don't know how to do it

OpenStudy (phi):

\[ \left(\sqrt[5]{x^7}\right)^3 \] replace the 5th root with the exponent 1/5 \[ \left((x^7)^\frac{1}{5}\right)^3 \]

OpenStudy (iwanttogotostanford):

ok

OpenStudy (iwanttogotostanford):

now what

OpenStudy (iwanttogotostanford):

? @phi

OpenStudy (phi):

work from the inside out using this rule \[ \left(a^b\right)^c = a^{bc} \]

OpenStudy (phi):

do you understand how to interpret that rule?